# UI-AGILE: Advancing GUI Agents with Effective Reinforcement Learning and Precise Inference-Time Grounding

### v3

@@ 4.1 Implementation Details
Data.
We collect data from multiple open-source datasets, including UI-R1 (Lu et al. 2025), GUI-R1 (Luo et al. 2025), Aguvis (Xu et al. 2024) and Grounding-R1 (Yang et al. 2025). We filter them using OmniParser (Wan et al. 2024) following Grounding-R1 (Yang et al. 2025). We randomly sample about 9k examples to train UI-AGILE.
Details of Environment, Implementation and Default Hyperparameters.
Due to space limit, we include the details of environment, implementation and default hyperparameters in Appendix A.
@@ 4.2 Grounding Capability Evaluation
Table 1: Grounding accuracy on ScreenSpot-Pro. “+Decomposed Grounding” denotes that the model uses our decomposed grounding with selection for enhancing inference. Results marked in bold represent the best performance, and underlined results indicate the second-best performance. The pass@4 metric indicates the success rate where a task is considered solved if the prediction of at least one of the sub-images is correct.
@@
Table 2: Grounding accuracy on ScreenSpot-v2. Results marked in bold represent the best performance, and underlined results indicate the second-best performance.
We evaluate the grounding ability on ScreenSpot-v2 (Wu et al. 2025) and ScreenSpot-Pro (Li et al. 2025). ScreenSpot-v2 is a corrected version of the original ScreenSpot (Cheng et al. 2024), providing evaluation of GUI grounding capability across mobile, desktop, and web platforms. ScreenSpot-Pro focuses on high-resolution professional environments, featuring expert-annotated tasks spanning 23 applications, five industries, and three operating systems. Since the images in ScreenSpot-v2 are already pre-cropped while those in ScreenSpot-Pro are full, uncropped displays, we evaluate our decomposed grounding with selection method exclusively on ScreenSpot-Pro.
Effectiveness of our inference enhancement.
As shown in Tab. 1, decomposed grounding with selection shows significant improvements on ScreenSpot-Pro. It provides a universal and substantial performance boost across all tested models, regardless of their original training paradigm (SFT or RFT). For instance, it elevates the average score of OS-Atlas-7B from 18.9 to 33.1 (+75.1%), and boosts Aguvis-7B from 20.4 to 36.5 (+78.9%). The consistent improvement proves the effectiveness of decomposed grounding with selection and its high applicability as a plug-and-play inference enhancement.
Effectiveness of our training enhancement.
Besides, Tab. 1 shows that UI-AGILE-3B and UI-AGILE-7B models, even without decomposed grounding, establish a new state-of-the-art among 3B and 7B models on ScreenSpot-Pro. Trained on only 9K examples with 2 epochs, they (37.9 for 3B and 44.0 for 7B) surpass other RFT-based models like UI-R1-E (33.5), InfiGUI-R1-3B and GUI-R1-7B (32.1). UI-AGILE-7B even outperforms the much larger model UI-TARS-72B (38.1) trained on approximately 50 billion tokens. On the ScreenSpot-v2 benchmark (Tab. 2), our UI-AGILE-7B also achieves state-of-the-art grounding accuracy with an average score of 92.1. The above results demonstrate the effectiveness of our proposed “Simple Thinking” reward, continuous grounding reward, and cropping-based resampling for improving the training of GUI agents.
@@ 4.3 Agent Capability Evaluation
Table 3: Type, GR and SR on AndroidControl-Low and AndroidControl-High. Results marked in bold represent the best performance, and underlined results indicate the second-best performance.
In addition to grounding-specific benchmarks, we also evaluate UI-AGILE-3B and UI-AGILE-7B on AndroidControl (Li et al. 2024) to assess its general agent capabilities.
Following the evaluation setting of OS-Atlas (Wu et al. 2025), we use three metrics: action type prediction accuracy (Type), grounding accuracy (GR), and the overall step success rate (SR). Type accuracy measures the exact match for the predicted action (e.g., click or scroll). For GR, a prediction is considered successful if it falls within a 14% screen-width radius of the ground-truth coordinate. The most holistic metric, SR, deems a step successful only if both the action type and all its associated arguments (e.g., coordinates for a click, direction for a scroll, or text for an input) are correct.
Following OS-Atlas, we use 7,708 examples for a fair comparison, while some works (e.g., Aguvis (Xu et al. 2024) and UGround (Gou et al. 2025)) randomly sample 500 action steps for testing. The evaluation is conducted under two distinct settings. In AndroidControl-Low, the agent receives a specific, low-level instruction for each step. In contrast, AndroidControl-High provides the agent with a high-level goal, requiring it to infer the correct action for the current step based on the conversation history and previous actions, thereby testing its multi-step reasoning capability.
As depicted in Tab. 3, UI-AGILE-7B achieves the best performance (SR: 77.6 and 60.6) compared to other RFT models, including UI-R1-E (SR: 71.37 and 35.88), GUI-R1-3B (SR: 64.41 and 46.55) and GUI-R1-7B (SR: 66.52 and 51.67). Notably, the smaller UI-AGILE-3B (SR: 77.4 and 56.9) also surpasses 7B baselines like GUI-R1-7B (SR: 66.5 and 51.7). For Type, a similar trend can be observed.
@@
Overall, the remarkable results on AndroidControl demonstrate that the improvements gained from our methods are not confined to improving grounding capability but also translate effectively to better decision-making and execution in complex, multi-step agent scenarios.
@@ 4.4 Ablation Study
Figure 3: Ablation study.
To verify the contribution of each training technique, we conduct an ablation study of UI-AGILE-3B and the results are shown in Fig. 3(a) and Fig. 3(b). Due to page limit, the performance on AndroidControl w.r.t. Type and GR are provided in Appendix C. We can observe that:
@@
  • Removing “Simple Thinking” during training (i.e., “No Thinking”) leads to higher grounding accuracy on ScreenSpot-Pro and ScreenSpot-v2 (0.4% and 0.7% gains). However, Fig. 3(b) shows that integrating “Simple Thinking” enhances the agent’s decision-making with planning, resulting in a noticeable improvement in SR on AndroidControl, with SR increases of 15.5% and 3.4% in Low and High settings, respectively.
4.5 Analysis of Inference Time
We report the inference time of our decomposed grounding with selection method on the full ScreenSpot-Pro dataset (Li et al. 2025) using the vLLM framework (Kwon et al. 2023) and one 80G A800 GPU card.
@@ 4.6 Analysis of Attempts Per Step
Due to space limit, we include the analysis of attempts per training step in Appendix D.
@@ 5 Conclusions
In this paper, we introduce UI-AGILE, a comprehensive framework designed to enhance GUI agents’ training and inference. It tackles the practical challenges of the reasoning-grounding dilemma, ineffective reward, and visual noise. Experimental results demonstrate the effectiveness of our proposed techniques on enhancing GUI agents’ grounding capability and general capability.
Despite the promising performance of UI-AGILE, the VLM used in its decomposed grounding with selection method is a general-purpose, pre-trained model that is not originally tuned for the selection task. A promising future direction would be to fine-tune this adjudicator model on a curated dataset of candidate UI elements, enhancing its selection accuracy to achieve further gains in overall grounding performance.
@@ Appendix A Environment and Implementation Details
Experiment Environment.
We use a machine with two Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz, 128GB main memory and eight NVIDIA A800 GPU for experiments.
Training Details.
We use the trl framework to implement the cropping-based resampling strategy and reward functions. The sampling process is attempted 4 times at most and is bypassed entirely if the bbox’s dimensions exceed the target crop size. Following prior works, we use Qwen2.5-VL-3B and Qwen2.5-VL-7B as base models.
Inference Details.
For decomposed grounding with selection, the input image is divided into four sub-images scaling to 60% of the original dimensions, with adjacent sub-images overlapping by 10% of the original image’s width and height. In the element image extraction stage, we define the element’s area by creating a simple bounding box centered on the predicted point with the width and height equal to 14% of the sub-image’s width and height. We have also explored a more sophisticated approach using OmniParser to refine this bounding box. However, it does not improve performance and increases the inference overhead. In the selection stage, we use Qwen2.5VL-7B-instruct to choose the final answer and the prompt is listed in Fig. 4.
Default Hyperparameters.
Tab. 4 provides the default hyperparameters of UI-AGILE where cropping factor is the width and height ratio of new attempted image and last attempted image.
@@
gradient accumulation steps 4
Table 4: Default Hyperparameters
Appendix B Code
@@
Appendix C Ablation Study
Fig 5 provides results of the ablation study using Type and GR as evaluation metrics on AndroidControl. We can observe that each component in UI-AGILE indeed contributes to the overall performance.
@@ Appendix C Ablation Study
Appendix D Analysis of Attempts Per Step
Fig. 6 shows the distribution of attempts per GRPO training step, where each step processes a batch of two training samples. In the first epoch, we find that only 61.8% of training steps are fully successful on the initial attempt (i.e., both samples in the batch are solved without resampling). This means that without our strategy, a minimum of 19.1% ($38.2\%\div 2$) of training samples would have provided no learning signal. Overall attempt numbers decreases in the second epoch, demonstrating that the model learns from the samples salvaged by our method.

### v4

@@ 4.1 Implementation Details
Data. We collect data from multiple open-source datasets, including UI-R1 [17], GUI-R1 [18], Aguvis [35] and Grounding-R1 [36]. We filter them using OmniParser [29] following Grounding-R1 [36]. We randomly sample 9k examples to train UI-AGILE. The value of hyperparameters in Eq. 2 and Eq. 5 are chosen by grid search.
Environment, Implementation and Hyperparameters. Due to space limit, we include the details in Appendix A and Appendix B.
@@ 4.2 Grounding Capability Evaluation
Table 1: Grounding accuracy on ScreenSpot-Pro. “+Decomposed Grounding” denotes that the model uses decomposed grounding with selection for enhancing inference. Bold and underlined Results represent the best performance and the second-best performance, respectively. pass@4 indicates the success rate where a task is considered solved if the prediction of at least one of the sub-images is correct.
@@
Table 2: Grounding accuracy on ScreenSpot-v2. Bold and underlined Results represent the best performance and the second-best performance, respectively.
@@
We evaluate the grounding ability on ScreenSpot-v2 [33] and ScreenSpot-Pro [12]. ScreenSpot-v2 is a corrected version of the original ScreenSpot [5], providing evaluation of GUI grounding capability across mobile, desktop, and web platforms. ScreenSpot-Pro focuses on high-resolution professional environments, featuring expert-annotated tasks spanning 23 applications, five industries, and three operating systems. Since the images in ScreenSpot-v2 are already pre-cropped while those in ScreenSpot-Pro are full, uncropped displays, we evaluate our decomposed grounding with selection method exclusively on ScreenSpot-Pro.
Effectiveness of Inference Enhancement. As shown in Tab. 1, decomposed grounding with selection shows significant improvements on ScreenSpot-Pro. It provides a universal and substantial performance boost across all tested models, regardless of their original training paradigm (SFT or RFT). For instance, it elevates the average score of OS-Atlas-7B from 18.9 to 33.1 (+75.1%), and boosts Aguvis-7B from 20.4 to 36.5 (+78.9%). The accuracy of selection (about 87%) can be inferred from pass@4 metric in Tab 1. The consistent improvement proves the effectiveness of decomposed grounding with selection and its high applicability as a plug-and-play inference enhancement.
Effectiveness of Training Enhancement. Besides, Tab. 1 shows that UI-AGILE-3B and UI-AGILE-7B models, even without decomposed grounding, establish a new state-of-the-art among 3B and 7B models on ScreenSpot-Pro. Trained on only 9K examples with 2 epochs, they (37.9 for 3B and 44.0 for 7B) surpass other RFT-based models like UI-R1-E (33.5), InfiGUI-R1-3B and GUI-R1-7B (32.1). UI-AGILE-7B even outperforms the much larger model UI-TARS-72B (38.1) trained on approximately 50 billion tokens. On the ScreenSpot-v2 benchmark (Tab. 2), our UI-AGILE-7B also achieves state-of-the-art grounding accuracy with an average score of 92.1. The above results demonstrate the effectiveness of our proposed “Simple Thinking” reward, continuous grounding reward, and cropping-based resampling for improving the training of GUI agents.
@@ 4.3 Agent Capability Evaluation
Table 3: Type, GR and SR on AndroidControl-Low and AndroidControl-High. Bold and underlined Results represent the best performance and the second-best performance, respectively.
@@
In addition to grounding-specific benchmarks, we also evaluate UI-AGILE-3B and UI-AGILE-7B on AndroidControl [13] to assess its general agent capabilities.
Following the evaluation setting of OS-Atlas [33], we use three metrics: action type prediction accuracy (Type), grounding accuracy (GR), and the overall step success rate (SR). Type accuracy measures the exact match for the predicted action (e.g., click or scroll). For GR, a prediction is considered successful if it falls within a 14% screen-width radius of the ground-truth coordinate. The SR deems a step successful only if both the action type and all its associated arguments (e.g., coordinates for a click, direction for a scroll, or text for an input) are correct.
Following OS-Atlas, we use 7,708 examples for a fair comparison, while some works (e.g., Aguvis [35] and UGround [7]) randomly sample 500 action steps for testing. The evaluation is conducted under two distinct settings. In AndroidControl-Low, the agent receives a specific, low-level instruction for each step. In contrast, AndroidControl-High provides the agent with a high-level goal, requiring it to infer the correct action for the current step based on the conversation history. We also note the strong performance of models like InfiGUI-R1 [15] and Aguvis [35] on AndroidControl. However, they leverage much larger datasets (32K/1M) compared to our 9K training samples, and InfiGUI-R1’s training data includes AndroidControl.
As depicted in Tab. 3, UI-AGILE-7B achieves the best performance (SR: 77.6 and 60.6) compared to other RFT models, including UI-R1-E (SR: 71.37 and 35.88), GUI-R1-3B (SR: 64.41 and 46.55) and GUI-R1-7B (SR: 66.52 and 51.67). For Type, a similar trend can be observed.
@@
Overall, the remarkable results on AndroidControl demonstrate that the improvements gained from our methods are not confined to improving grounding capability but also translate effectively to better decision-making in multi-step agent scenarios.
@@ 4.4 Ablation Study
Figure 3: Ablation result on the three benchmarks.
Ablation Experiments on Training. To verify the contribution of each training technique, we conduct an ablation study of UI-AGILE-3B and the results are shown in Fig. 3(a) and Fig. 3(b). Due to page limit, the performance on AndroidControl w.r.t. Type and GR are provided in Appendix C. From the results, we can observe that:
@@
  • Comparing the three strategies (“No Thinking”, “Simple Thinking”, “Normal Thinking”), we can observe that:
    • Grounding accuracy descends, while Android Control accuracy ascends.
    • “No Thinking” performs significantly worse on the Android Control task.
    • “Simple Thinking” is more accurate in grounding than “Normal Thinking” yet achieves comparable performance on Android Control.
    • “Normal Thinking” requires 1.7x more training time than “Simple Thinking”.
    Hence, we can conclude that “Simple Thinking” can achieve the balance between effectiveness and efficiency.
Analysis of Selection Accuracy For our Decomposed Grounding with Selection method, the selection accuracy, inferred from the pass@4 metric in Tab. 1, is 82.26%. We attempted to fine-tune the selection VLM, teaching the model to output “Yes” if the intruction matches the element otherwise “No”. The hard negative samples, which are instruction-element pairs, are genetated from the grounding model’s own predictions when using our Decomposed Grounding with Selection method. However, this fine-tuning dropped selection accuracy to 68%, likely due to a mismatch between the training task and the score-based ranking task during inference.
We then explored an alternative: training the VLM to output the index of the correct element image from four candidates given the instruction. This approach perfectly aligns the training and inference tasks. Training on 9k data (identical to the dataset used to train UI-AGILE) improved selection accuracy from  70% to 81%.
@@ 4.5 Analysis of Attempts Per Step
Fig. 4 shows the distribution of attempts per GRPO training step using cropping-based resampling strategy, where each step processes a batch of two training samples. In the first epoch, we find that only 61.8% of training steps are fully successful on the initial attempt (i.e., both samples in the batch are solved without resampling). This means that without our strategy, 19.1% to 38.2% of training samples would have provided no learning signal. Overall attempt numbers decrease in the second epoch, demonstrating that the model learns from the samples salvaged by our cropping-based resampling.
5 Related Work
Reinforcement Learning for Large Models. Recently, RL algorithm, including PPO [25], DPO [24] and GRPO [26], for training large models has gained significant momentum. These algorithms have achieved remarkable success in enhancing the reasoning capabilities of large models on complex tasks, with models like OpenAI O1 [9] and DeepSeek-R1 [6]. The efficacy of these approaches promotes their rapid extension into the multimodal domain [3, 4, 16, 22, 20, 19, 32].
GUI Agents. Techniques for GUI agents have evolved rapidly [30, 11] in recent years. Following early works like CogAgent [8] and SeeClick [5], most studies trained models to output actions (e.g., clicking with coordinate parameters, keyboard input) based on visual inputs screenshots and user instructions. Show-UI [14] innovates on visual processing efficiency. OS-Atlas [33], UGround [7] and Aria-UI [37] propose novel, large-scale pipelines to collect and synthesize millions of GUI agent trajectories, significantly improving model generalization. Aguvis [35] introduced a two-stage training process that explicitly uses VLM-generated Chain-of-Thought (CoT) data to teach planning and reasoning. JEDI [34] constructs a refusal part by mismatching existing instructions with unrelated screenshots. Standing out in complexity and scale, UI-TARS [23] utilizes the largest dataset and the most intricate training pipeline, which involves SFT and DPO on human annotated CoT data to improve performance. This data-intensive scaling has motivated a shift towards more effective Reinforcement Learning with Verifiable Rewards paradigms, first explored by UI-R1 [17] and GUI-R1 [18]. InfiGUI-R1 [15] employs Spatial Reasoning Distillation to enhance cross-modal spatial reasoning capabilities and uses RL to refine the basic reasoner into a deliberative one. GUI-G1 [39] leverages Hit-based reward and IoU-based reward for improving GUI agents. RULER [31] uses tokens as explicit coordinate markers, letting the model reference positions similar to gridlines.
@@ 6 Conclusions
In this paper, we introduce UI-AGILE, a comprehensive framework designed to enhance GUI agents’ training and inference. It tackles the practical challenges of the reasoning-grounding dilemma, ineffective reward, and visual noise. Experimental results demonstrate the effectiveness of our proposed techniques on enhancing GUI agents. Future work may include exploring alternative selection stage methods, such as computing text-element image embedding similarity and cropping as a function call.
7 Acknowledgment
This work is supported by the National Key Research and Development Program of China (No. 2025YFE0113500), the National Science Fund for Distinguished Young Scholars (No. 62525605), and the National Natural Science Foundation of China (No. 62572410, No. 62272401, and No. U22B2051).
Supplementary Material
@@ Appendix A Environment and Implementation Details
Environment. We use a machine with two Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz, 512GB main memory and eight NVIDIA A800 GPU for experiments.
Training Details. We use the trl framework [^1] https://github.com/huggingface/trl to implement the cropping-based resampling strategy and reward functions. The sampling process is attempted 4 times at most and is bypassed entirely if the bbox’s dimensions exceed the target crop size. Following prior works, we use Qwen2.5-VL-3B [^2] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B [^3] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct as base models.
Inference Details. For decomposed grounding with selection, the input image is divided into four sub-images scaling to 60% of the original dimensions, with adjacent sub-images overlapping by 10% of the original image’s width and height. In the element image extraction stage, we define the element’s area by creating a simple bounding box centered on the predicted point with the width and height equal to 14% of the sub-image’s width and height. We have also explored a more sophisticated approach using OmniParser to refine this bounding box. However, it does not improve performance and increases the inference overhead. In the selection stage, we use Qwen2.5VL-7B-instruct to choose the final answer and the prompt is listed in Fig. 5.
Hyperparameters. Tab. 4 provides the training hyperparameters of UI-AGILE where cropping factor is the width and height ratio of new attempted image and last attempted image. The hyperparameters in Eq. 2 and Eq. 5 are optimized using a grid search. Tab. 5 shows the specific values.
Appendix B Code
Table 4: Training hyperparameters.
@@ Appendix B Code
sampling attempt num 4
Table 5: Hyperparameters for Eq. 2 and Eq. 5
Hyperparameter Value
$l_{\text{ideal\_start}}$ 120 chars
$l_{\text{ideal\_end}}$ 200 chars
$l_{\text{min}}$ 50 chars
$l_{\text{max}}$ 300 chars
$C$ 4
@@
Appendix C Additional Results for Ablation Study
Fig. 6 provides results of the ablation study using Type and GR as evaluation metrics on AndroidControl. We can observe that each component in UI-AGILE indeed contributes to the overall performance.
@@ Appendix C Additional Results for Ablation Study
Appendix D Analysis of Inference Time
We report the inference time of our decomposed grounding with selection method on the full ScreenSpot-Pro dataset [12] using the vLLM framework [10] and one 80G A800 GPU card.